OpenStudy (anonymous):

Explain the meaning of absolute value of a complex number. What is the absolute value of –2 + 3i?

OpenStudy (anonymous):

The absolute value of a+bi is sqrt(a^2 + b^2)

OpenStudy (solomonzelman):

absolute value means, lets take a number line,

OpenStudy (anonymous):

So the absolute-value of 3-4i is sqrt(3^2 + (-4)^2) = sqrt(9 + 16) = sqrt (25) = 5.

OpenStudy (solomonzelman):

For complex number like –2 + 3i it will be –2 + 3i positive or –2 + 3i negative. In toher words +2-3i or -2+3i

OpenStudy (anonymous):

so the first part?